OpenStudy (daniellelovee):

Calculate the slope of the line in your graph of the square of the period of the pendulum vs. length of the string [slope = (y2 – y1)/(x2 – x1)]. Galileo figured out the equation that describes the behavior of a pendulum. If you square both sides of the equation, you will find that the slope of the line is related to the acceleration due to gravity (g). Specifically, slope = 42/g. Use your data to calculate g. How does it compare with the accepted value of 9.807 m/s2? @satellite73

OpenStudy (daniellelovee):

@Astrophysics

OpenStudy (astrophysics):

What is the equation, and what does your graph represent?

OpenStudy (daniellelovee):

g = 4π² / slope.

OpenStudy (daniellelovee):

my slope?

OpenStudy (astrophysics):

No I'm asking, the formula you're testing is \[T =2 \pi \sqrt{\frac{ L }{ g }}\] correct

OpenStudy (daniellelovee):

oh ok yeah and I can also use T² = 4π²L/g

OpenStudy (astrophysics):

Ok and what does your graph represent, (y as a function of x) state the variables

OpenStudy (daniellelovee):

thats when I get lost sorry

OpenStudy (astrophysics):

Ok no worries, note that we are determining gravity right

OpenStudy (daniellelovee):

yes

OpenStudy (astrophysics):

Ok so you gathered that data right in the table?

OpenStudy (daniellelovee):

yes

OpenStudy (astrophysics):

Ok cool, so what do the table represent?

OpenStudy (astrophysics):